OpenStudy (anonymous):

Simplify the expression. (15 + 8)[23 – (10 + 7)]

OpenStudy (kewlgeek555):

Well, let's follow PEMDAS. Parenthesis Exponents Multiplication or Division (left to right) Addition OR Subtraction (left to right)

OpenStudy (kewlgeek555):

Which expression do you think we need to solve first?

OpenStudy (anonymous):

138

OpenStudy (anonymous):

parentesis

OpenStudy (anonymous):

its not 138 sorry

OpenStudy (kewlgeek555):

@crabby_patty : Please read the code of conduct, as this is a study site, not a "give an answer" website. Thank you.

OpenStudy (anonymous):

fine

OpenStudy (kewlgeek555):

And @mimi723 : But which equation inside of a parenthesis do we solve first? (15 + 8) [23 - (10 + 7)] OR (10 + 7) HINT: Parenthesis inside of Parenthesis come first. c;

OpenStudy (anonymous):

oops i mean (10+7)

OpenStudy (kewlgeek555):

Yes, the first one is (15 + 8), but it would be safer to solve (10 + 7) first, since it is inside of another set of parenthesis. So... \[(15+8)[23-(10+7)]\]\[(15+8)[23-17]\]Do you know how to solve it from here?

OpenStudy (anonymous):

oops i typed that wrong

OpenStudy (anonymous):

no because i would just have 6 and 23 and have nothing to do with them

OpenStudy (kewlgeek555):

Hm. I really do not understand what you are saying there. I also want to ask, is it.. (15 + 8)[23 - 17] OR (15 + 8) - [23 - 17]

OpenStudy (anonymous):

(15 + 8)[23 - 17]

OpenStudy (kewlgeek555):

Oh, well if it is like that, then it means that it is 23 TIMES 6. When two parenthesis are together, they are being multiplied... \[(15+8)[23-17]=(15+8)\times[23-17]=23\times6=???\] So, do you know what 23 times 6 is?

OpenStudy (anonymous):

not by heart

OpenStudy (anonymous):

138?

OpenStudy (kewlgeek555):

Yes.

OpenStudy (anonymous):

yay

OpenStudy (anonymous):

thnx a lot

OpenStudy (kewlgeek555):

You are welcome. I will give you a medal for being such a great mathematician. ;D

OpenStudy (kewlgeek555):

Thanks for the fan as well.

OpenStudy (anonymous):

thank you :D